5a-2 over3 equals -9
steps to get answer so I understand how to do it







Answer by JulietG(1812)

(5a-2)/3 = -9
Multiply each side by 3 to get rid of the fraction
5a-2 = -9 * 3
5a-2 = -27
Add 2 to each side
5a = -25
Divide each side by 5
a = -5